# MatDecompSDF: High-Fidelity 3D Shape and PBR Material Decomposition from Multi-View Images

### v2

@@ 1. Introduction
The demand for high-fidelity, interactive, and editable 3D content has become a cornerstone of modern digital experiences, fueling advancements in virtual and augmented reality, digital twins, and the entertainment industry . Traditionally, creating such photorealistic digital assets is a labor-intensive process, often requiring expert artists to manually model geometry and define complex material properties. The advent of neural scene representations has marked a paradigm shift, offering a new avenue for synthesizing and reconstructing the 3D world directly from images. Groundbreaking work on Neural Radiance Fields (NeRF) demonstrated an unprecedented ability to synthesize photorealistic novel views from a collection of images by representing a scene as a continuous volumetric function (mildenhall2020nerf). This seminal work spurred a wave of research exploring dynamic scenes (pumarola2021dnerf; li2021realtimedynamic) and even articulated subjects like humans (li2023tava; wu2023magicpony). However, the implicit volumetric nature of NeRF makes it challenging to extract explicit, high-quality surface geometry, which is crucial for most graphics pipelines.
@@
Beyond the visual spectrum, the broader field of sensing is also moving towards extracting nuanced information from ubiquitous signals. The ability to recognize human activities (huang2023phaseanti), gestures (gu2023attentionbased), emotions (gu2023wife), and even vital pulmonary functions (zhao2025wipulmo; wang2024targetoriented) from commodity WiFi signals showcases a parallel trend of disentangling complex information from noisy, indirect measurements. The security and privacy implications of such advanced sensing, from fingerprinting attacks (huang2024phyfinatt) to keystroke eavesdropping (huang2024keystrokesniffer), highlight a universal need for robust and secure systems, a consideration that will become paramount as 3D reconstruction moves towards distributed, user-sourced data. Architectures for such distributed systems, potentially leveraging federated learning with neural architecture search (liu2024finch) and experience-driven model migration (liu2024federated), provide a roadmap for future privacy-preserving 3D capture ecosystems. This thesis situates itself within this rich context, aiming to push the boundaries of high-fidelity, physically-based inverse rendering of general objects. We focus on robustly and efficiently disentangling geometry and complex material properties from multi-view images, drawing inspiration from both the explicit structure of physically-based rendering and the powerful priors of modern neural networks (wu2024extracting).
@@ 2.2. Neural Inverse Rendering
The current state-of-the-art is represented by powerful frameworks like NVDIFFREC (munkberg2022nvdiffrec), which directly optimizes a triangle mesh, its PBR materials, and an HDR light map to match input images. Mat-NeuS (zhang2022matneus) extends the high-fidelity surface reconstruction of NeuS to simultaneously acquire material properties. Neural-PIL (boss2022neuralpil) introduced pre-integrated lighting to efficiently handle complex reflectance. These methods have been successfully applied to reconstruct general objects (zhang2023invrender) and even full indoor scenes (yen2021nerfilab). The core challenge in all these methods is robust disentanglement, which requires careful regularization and priors. This is conceptually similar to the need for robust feature extraction in other domains, such as developing anti-interference techniques for WiFi-based activity recognition (huang2023phaseanti) or building secure authentication systems resilient to fingerprinting attacks (huang2024phyfinatt). Our work contributes to this line of research by proposing new techniques to improve the robustness of material decomposition under unknown lighting. We also draw inspiration from how hierarchical models can parse complex structures, a principle proven effective in tasks like sign language translation.
@@ 2.3. Generative Priors for 3D Content Creation
The rise of large-scale generative models, particularly 2D diffusion models, has unlocked new paradigms for 3D content creation. DreamFusion (poole2022dreamfusion) introduced Score Distillation Sampling (SDS), a method to ”distill” the knowledge of a pre-trained 2D text-to-image model to guide the optimization of a 3D representation like NeRF. This spawned a flurry of research in text-to-3D generation, with follow-up works like Magic3D (lin2023magic3d) improving resolution and efficiency. ProlificDreamer (wang2023prolificdreamer) proposed Variational Score Distillation (VSD) to enhance the diversity and fidelity of generated results. A key focus of recent work is the disentanglement of geometry and appearance within this generative framework. Fantasia3D (chen2023fantasia3d) and TexFusion (seo2023texfusion) explicitly aim to generate not just a shape, but also assign meaningful textures and PBR materials based on textual prompts. This leverages the powerful multi-modal understanding of joint vision-language models like CLIP. Such models are also used for text-guided editing of existing 3D shapes (richardson2023texture) or neural fields (wang2022clipnerf), enabling intuitive manipulation of scene properties (dadashnialehi2023neuralscenedecoration). The ability to generate content based on abstract, multi-modal inputs resonates with research in visual dialog , video grounding , and even cross-modal emotion recognition (gu2023wife), all of which seek to bridge the gap between different data modalities.
@@ 2.4. Modeling of Advanced Materials and Dynamic Scenes
Modeling dynamic and deformable objects presents another set of challenges. D-NeRF (pumarola2021dnerf) pioneered the use of a deformation field to map a dynamic scene to a canonical space. This concept has been extended to create controllable and animatable avatars of human actors , often incorporating relighting capabilities (zhao2022relightableavatar). The ability to track and model human motion finds parallels in other sensing modalities, such as gesture recognition using WiFi (gu2023attentionbased). The overarching goal is to build robust models from diverse, often unstructured data, whether it’s from online image collections (liu2022neroic) or from distributed agents in a federated learning setting (liu2024finch; liu2024federated). The security and privacy of such data are paramount, a concern highlighted by research into the vulnerabilities of sensing systems (huang2024keystrokesniffer). Our work focuses on static objects but aims for a material representation robust enough to serve as a foundation for future extensions into these more complex dynamic and material domains, where understanding subtle signals, much like in audio-visual event analysis or respiratory healthcare sensing (zhao2025wipulmo; wang2024targetoriented), is key.
@@ 3.4. Optimization and Loss Functions
The networks are trained end-to-end by minimizing a composite loss function. The ability to learn from complex, multi-modal, and sometimes noisy data is crucial, a challenge shared by fields as diverse as sign language production and WiFi-based sensing (wang2024targetoriented).
@@ 3.4.2. Geometric Regularization
We also use a mask loss to enforce that SDF values inside the object (as defined by the input masks) are negative. This robust supervision is critical, much like how robust authentication is needed to prevent attacks in other sensing domains (huang2024phyfinatt).
@@ 3.4.3. Material and Lighting Priors
Additionally, we add a sparsity prior on the metallic parameter to encourage it to be close to either 0 (dielectric) or 1 (metal), reflecting the behavior of most common materials. We also regularize the total intensity and smoothness of the predicted environment light map from $f_{light}$ to prevent overly complex lighting from being ”baked” into the material albedo. The design of these priors is crucial for stable training, a principle that holds true when designing complex hierarchical systems, whether for federated learning (liu2024finch) or visual understanding .
@@ 4.1.1. Datasets
We evaluate our method on both synthetic and real-world datasets to demonstrate its effectiveness and generalization capabilities. The diversity of data is crucial for robust evaluation, a principle that is vital across various domains, from training models for sign language to ensuring robust performance in WiFi-based activity recognition under interference (huang2023phaseanti).
@@
To validate our method on real-world data, we use the widely-adopted DTU dataset . The DTU dataset contains multi-view images of various objects captured under controlled laboratory conditions with a fixed camera rig and several different lighting conditions. This dataset is a standard benchmark for multi-view surface reconstruction (wang2021neus; yariv2021volsdf). Since ground-truth material properties are not available for these real objects, evaluation is primarily focused on the quality of the geometric reconstruction and the visual fidelity of novel view synthesis and relighting results. The challenge of working with real-world data, where ground truth is scarce, mirrors challenges in other sensing domains that rely on unsupervised or weakly-supervised methods, such as gesture recognition from commodity signals (gu2023attentionbased).
@@ 4.1.2. Evaluation Metrics
  • Novel View Synthesis: For both datasets, we evaluate the quality of rendered novel views against ground-truth images using Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index (SSIM), and the Learned Perceptual Image Patch Similarity (LPIPS) (srinivasan2023nero). PSNR and SSIM measure pixel-level fidelity, while LPIPS better correlates with human perceptual judgment, a crucial aspect for tasks aiming at human-centric understanding, such as visual dialog or emotion recognition (gu2023wife).
@@ 4.1.3. Baselines and Implementation Details
Our framework is implemented in PyTorch. All experiments are conducted on NVIDIA A100 GPUs with 40GB of memory. We use the Adam optimizer with an initial learning rate of $5\times 10^{-4}$, which decays exponentially over the course of training. Each model is trained for 300k iterations with a batch size of 512 rays. The loss weights are set empirically, with $\lambda_{L1}=1.0$, $\lambda_{LPIPS}=0.2$, $\lambda_{eikonal}=0.1$, and $\lambda_{mat}=0.01$. The design of an efficient and stable training pipeline is paramount, a concern shared by those developing complex distributed systems for tasks like federated learning (liu2024finch; liu2024federated) or dealing with the security implications of advanced sensing technologies (huang2024keystrokesniffer). Our code and pre-trained models will be made publicly available to facilitate future research.
@@ 4.2. Quantitative Comparison
Table 2 shows the performance on the challenging real-world DTU dataset. Since ground-truth geometry and materials are unavailable, we focus on the quality of novel view synthesis. Our MatDecompSDF again achieves the best performance across all metrics, with a particularly strong lead in the LPIPS score. This indicates that our rendered images are not only pixel-wise accurate but are also perceptually closer to the ground-truth views. We attribute this to two factors: first, the inclusion of the $\mathcal{L}_{LPIPS}$ term in our photometric loss guides the optimization towards perceptually plausible solutions, a strategy whose importance is recognized in generative modeling (srinivasan2023nero). Second, by modeling the scene with physically accurate materials and lighting, our framework produces more realistic shading and highlights, which are key components of perceptual quality. This is in contrast to geometry-only methods like NeuS, whose view-dependent appearance model can sometimes produce results that look flat or lack realism under complex lighting. The ability to perform robustly on real-world data with its inherent noise and complexities is a testament to our framework’s design, echoing the need for robust systems in other real-world sensing applications, from federated learning in heterogeneous networks (liu2024federated) to secure WiFi-based gesture recognition (gu2023attentionbased).
@@ 4.3. Qualitative Comparison
Figure 2 visualizes the decomposed PBR material maps for an object from our synthetic dataset, compared against leading baselines. The visual difference is striking. The albedo map produced by our MatDecompSDF is significantly cleaner, showing pure surface colors without the shadowy artifacts present in the results of NVDIFFREC (munkberg2022nvdiffrec) and Mat-NeuS (zhang2022matneus). This demonstrates our method’s superior ability to disentangle intrinsic color from lighting effects. Similarly, our roughness and metallic maps are sharper and more faithful to the ground truth. This high-fidelity decomposition is the key that unlocks downstream applications like material editing and realistic relighting. The clarity of our decomposition can be likened to the goal of achieving interference-independent feature extraction in WiFi sensing (huang2023phaseanti).
@@ 4.5. Discussion and Limitations
Finally, as 3D capture technology becomes more democratized, it will increasingly rely on data from distributed, personal devices. This raises important questions about privacy and security. The ability to reconstruct detailed 3D models could be misused, and the underlying sensing systems could be vulnerable, a concern echoed in the domain of WiFi sensing which has demonstrated risks like keystroke sniffing (huang2024keystrokesniffer). Future work should explore privacy-preserving reconstruction techniques, potentially leveraging federated learning frameworks (liu2024federated) to train models without centralizing sensitive user data, and ensuring the security of the underlying capture and processing pipeline.

### v3

@@ 1. Introduction
The demand for high-fidelity, interactive, and editable 3D content has become a cornerstone of modern digital experiences, fueling advancements in virtual and augmented reality, digital twins, and the entertainment industry . Traditionally, creating such photorealistic digital assets is a labor-intensive process, often requiring expert artists to manually model geometry and define complex material properties. The advent of neural scene representations has marked a paradigm shift, offering a new avenue for synthesizing and reconstructing the 3D world directly from images. Groundbreaking work on Neural Radiance Fields (NeRF) demonstrated an unprecedented ability to synthesize photorealistic novel views from a collection of images by representing a scene as a continuous volumetric function (mildenhall2020nerf). This seminal work spurred a wave of research exploring dynamic scenes (pumarola2021dnerf) and even articulated subjects like humans (li2023tava; wu2023magicpony). However, the implicit volumetric nature of NeRF makes it challenging to extract explicit, high-quality surface geometry, which is crucial for most graphics pipelines.
@@
Beyond the visual spectrum, the broader field of sensing is also moving towards extracting nuanced information from ubiquitous signals. The ability to recognize human activities , gestures , emotions , and even vital pulmonary functions from commodity WiFi signals showcases a parallel trend of disentangling complex information from noisy, indirect measurements. The security and privacy implications of such advanced sensing, from fingerprinting attacks to keystroke eavesdropping , highlight a universal need for robust and secure systems, a consideration that will become paramount as 3D reconstruction moves towards distributed, user-sourced data. Architectures for such distributed systems, potentially leveraging federated learning with neural architecture search and experience-driven model migration , provide a roadmap for future privacy-preserving 3D capture ecosystems. This thesis situates itself within this rich context, aiming to push the boundaries of high-fidelity, physically-based inverse rendering of general objects. We focus on robustly and efficiently disentangling geometry and complex material properties from multi-view images, drawing inspiration from both the explicit structure of physically-based rendering and the powerful priors of modern neural networks (wu2024extracting).
@@ 2.2. Neural Inverse Rendering
The current state-of-the-art is represented by powerful frameworks like NVDIFFREC (munkberg2022nvdiffrec), which directly optimizes a triangle mesh, its PBR materials, and an HDR light map to match input images. Mat-NeuS (zhang2022matneus) extends the high-fidelity surface reconstruction of NeuS to simultaneously acquire material properties. Neural-PIL (boss2022neuralpil) introduced pre-integrated lighting to efficiently handle complex reflectance. These methods have been successfully applied to reconstruct general objects (zhang2023invrender) and even full indoor scenes (yen2021nerfilab). The core challenge in all these methods is robust disentanglement, which requires careful regularization and priors. This is conceptually similar to the need for robust feature extraction in other domains, such as developing anti-interference techniques for WiFi-based activity recognition or building secure authentication systems resilient to fingerprinting attacks . Our work contributes to this line of research by proposing new techniques to improve the robustness of material decomposition under unknown lighting. We also draw inspiration from how hierarchical models can parse complex structures, a principle proven effective in tasks like sign language translation.
@@ 2.3. Generative Priors for 3D Content Creation
The rise of large-scale generative models, particularly 2D diffusion models, has unlocked new paradigms for 3D content creation. DreamFusion (poole2022dreamfusion) introduced Score Distillation Sampling (SDS), a method to ”distill” the knowledge of a pre-trained 2D text-to-image model to guide the optimization of a 3D representation like NeRF. This spawned a flurry of research in text-to-3D generation, with follow-up works like Magic3D (lin2023magic3d) improving resolution and efficiency. ProlificDreamer (wang2023prolificdreamer) proposed Variational Score Distillation (VSD) to enhance the diversity and fidelity of generated results. A key focus of recent work is the disentanglement of geometry and appearance within this generative framework. Fantasia3D (chen2023fantasia3d) and TexFusion (seo2023texfusion) explicitly aim to generate not just a shape, but also assign meaningful textures and PBR materials based on textual prompts. This leverages the powerful multi-modal understanding of joint vision-language models like CLIP. Such models are also used for text-guided editing of existing 3D shapes (richardson2023texture) or neural fields (wang2022clipnerf), enabling intuitive manipulation of scene properties (dadashnialehi2023neuralscenedecoration). The ability to generate content based on abstract, multi-modal inputs resonates with research in visual dialog , video grounding , and even cross-modal emotion recognition , all of which seek to bridge the gap between different data modalities.
@@ 2.4. Modeling of Advanced Materials and Dynamic Scenes
Modeling dynamic and deformable objects presents another set of challenges. D-NeRF (pumarola2021dnerf) pioneered the use of a deformation field to map a dynamic scene to a canonical space. This concept has been extended to create controllable and animatable avatars of human actors , often incorporating relighting capabilities (zhao2022relightableavatar). The ability to track and model human motion finds parallels in other sensing modalities, such as gesture recognition using WiFi . The overarching goal is to build robust models from diverse, often unstructured data, whether it’s from online image collections (liu2022neroic) or from distributed agents in a federated learning setting . The security and privacy of such data are paramount, a concern highlighted by research into the vulnerabilities of sensing systems . Our work focuses on static objects but aims for a material representation robust enough to serve as a foundation for future extensions into these more complex dynamic and material domains, where understanding subtle signals, much like in audio-visual event analysis or respiratory healthcare sensing , is key.
@@ 3.4. Optimization and Loss Functions
The networks are trained end-to-end by minimizing a composite loss function. The ability to learn from complex, multi-modal, and sometimes noisy data is crucial, a challenge shared by fields as diverse as sign language production and WiFi-based sensing .
@@ 3.4.2. Geometric Regularization
We also use a mask loss to enforce that SDF values inside the object (as defined by the input masks) are negative. This robust supervision is critical, much like how robust authentication is needed to prevent attacks in other sensing domains .
@@ 3.4.3. Material and Lighting Priors
Additionally, we add a sparsity prior on the metallic parameter to encourage it to be close to either 0 (dielectric) or 1 (metal), reflecting the behavior of most common materials. We also regularize the total intensity and smoothness of the predicted environment light map from $f_{light}$ to prevent overly complex lighting from being ”baked” into the material albedo. The design of these priors is crucial for stable training, a principle that holds true when designing complex hierarchical systems, whether for federated learning or visual understanding .
@@ 4.1.1. Datasets
We evaluate our method on both synthetic and real-world datasets to demonstrate its effectiveness and generalization capabilities. The diversity of data is crucial for robust evaluation, a principle that is vital across various domains, from training models for sign language to ensuring robust performance in WiFi-based activity recognition under interference.
@@
To validate our method on real-world data, we use the widely-adopted DTU dataset . The DTU dataset contains multi-view images of various objects captured under controlled laboratory conditions with a fixed camera rig and several different lighting conditions. This dataset is a standard benchmark for multi-view surface reconstruction (wang2021neus; yariv2021volsdf). Since ground-truth material properties are not available for these real objects, evaluation is primarily focused on the quality of the geometric reconstruction and the visual fidelity of novel view synthesis and relighting results. The challenge of working with real-world data, where ground truth is scarce, mirrors challenges in other sensing domains that rely on unsupervised or weakly-supervised methods, such as gesture recognition from commodity signals .
@@ 4.1.2. Evaluation Metrics
  • Novel View Synthesis: For both datasets, we evaluate the quality of rendered novel views against ground-truth images using Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index (SSIM), and the Learned Perceptual Image Patch Similarity (LPIPS) (srinivasan2023nero). PSNR and SSIM measure pixel-level fidelity, while LPIPS better correlates with human perceptual judgment, a crucial aspect for tasks aiming at human-centric understanding, such as visual dialog or emotion recognition .
@@ 4.1.3. Baselines and Implementation Details
Our framework is implemented in PyTorch. All experiments are conducted on NVIDIA A100 GPUs with 40GB of memory. We use the Adam optimizer with an initial learning rate of $5\times 10^{-4}$, which decays exponentially over the course of training. Each model is trained for 300k iterations with a batch size of 512 rays. The loss weights are set empirically, with $\lambda_{L1}=1.0$, $\lambda_{LPIPS}=0.2$, $\lambda_{eikonal}=0.1$, and $\lambda_{mat}=0.01$. The design of an efficient and stable training pipeline is paramount, a concern shared by those developing complex distributed systems for tasks like federated learning or dealing with the security implications of advanced sensing technologies . Our code and pre-trained models will be made publicly available to facilitate future research.
@@ 4.2. Quantitative Comparison
Table 2 shows the performance on the challenging real-world DTU dataset. Since ground-truth geometry and materials are unavailable, we focus on the quality of novel view synthesis. Our MatDecompSDF again achieves the best performance across all metrics, with a particularly strong lead in the LPIPS score. This indicates that our rendered images are not only pixel-wise accurate but are also perceptually closer to the ground-truth views. We attribute this to two factors: first, the inclusion of the $\mathcal{L}_{LPIPS}$ term in our photometric loss guides the optimization towards perceptually plausible solutions, a strategy whose importance is recognized in generative modeling (srinivasan2023nero). Second, by modeling the scene with physically accurate materials and lighting, our framework produces more realistic shading and highlights, which are key components of perceptual quality. This is in contrast to geometry-only methods like NeuS, whose view-dependent appearance model can sometimes produce results that look flat or lack realism under complex lighting. The ability to perform robustly on real-world data with its inherent noise and complexities is a testament to our framework’s design, echoing the need for robust systems in other real-world sensing applications, from federated learning in heterogeneous networks
@@ 4.3. Qualitative Comparison
Figure 2 visualizes the decomposed PBR material maps for an object from our synthetic dataset, compared against leading baselines. The visual difference is striking. The albedo map produced by our MatDecompSDF is significantly cleaner, showing pure surface colors without the shadowy artifacts present in the results of NVDIFFREC (munkberg2022nvdiffrec) and Mat-NeuS (zhang2022matneus). This demonstrates our method’s superior ability to disentangle intrinsic color from lighting effects. Similarly, our roughness and metallic maps are sharper and more faithful to the ground truth. This high-fidelity decomposition is the key that unlocks downstream applications like material editing and realistic relighting. The clarity of our decomposition can be likened to the goal of achieving interference-independent feature extraction in WiFi sensing.
@@ 4.5. Discussion and Limitations
Finally, as 3D capture technology becomes more democratized, it will increasingly rely on data from distributed, personal devices. This raises important questions about privacy and security. The ability to reconstruct detailed 3D models could be misused, and the underlying sensing systems could be vulnerable, a concern echoed in the domain of WiFi sensing which has demonstrated risks like keystroke sniffing . Future work should explore privacy-preserving reconstruction techniques, potentially leveraging federated learning frameworks to train models without centralizing sensitive user data, and ensuring the security of the underlying capture and processing pipeline.